# CC-GPX: Extracting High-Quality Annotated Geospatial Data from Common Crawl

### v2

@@ Abstract.
The Common Crawl (CC) corpus is the largest open web crawl dataset containing 9.5+ petabytes of data captured since 2008. The dataset is instrumental in training large language models, and as such it has been studied for (un)desirable content, and distilled for smaller, domain-specific datasets. However, to our knowledge, no research has been dedicated to using CC as a source of annotated geospatial data. In this paper, we introduce an efficient pipeline to extract annotated user-generated tracks from GPX files found in CC, and the resulting multimodal dataset with 1,416 pairings of human-written descriptions and MultiLineString vector data from the 6 most recent CC releases. The dataset can be used to study people’s outdoor activity patterns, the way people talk about their outdoor experiences, and for developing trajectory generation or track annotation models. Our reproducible code is available on GitHub [^1] https://github.com/ilyankou/cc-gpx.
@@ 2.5. Translating descriptions to English
The majority of GPX files we identified are from European websites, and their descriptions are not in English. We use pycld2131313https://github.com/aboSamoor/pycld2 Python library to identify the original language. We remove tracks whose language is identified as Unknown as those often contain majority numbers or incoherent text, as well as all tracks whose descriptions are in languages that have 5 or fewer samples.
@@ 3. Dataset
4. Applications
The tracks in the resulting dataset represent real-life outdoor activities that were either recorded (i.e., completed) or planned using GIS software. Thus, these routes can be used in training of GPS trajectory generation models that are focused on recreational activities.
The description-MultiLineString pairs can be used to study how people describe their outdoor activities, such as what kinds of landmarks, landscapes, and experiences they choose to mention or ignore, and get an insight into turn-by-turn decisions of users. The immediate application can be to fine-tine language models to generate human-like descriptions of existing or synthetically composed routes. This dataset can become especially powerful when combined with Point of Interest (POI) and/or road segment datasets, such as those from OpenStreetMap [^18] https://welcome.openstreetmap.org/ or Overture [^19] https://overturemaps.org/.
@@ Acknowledgements.
This work was supported by the Ordnance Survey, and the Engineering and Physical Sciences Research Council [grant no. EP/Y528651/1].

### v3

@@ Abstract.
The Common Crawl (CC) corpus is the largest open web crawl dataset containing 9.5+ petabytes of data captured since 2008. The dataset is instrumental in training large language models, and as such it has been studied for (un)desirable content, and distilled for smaller, domain-specific datasets. However, to our knowledge, no research has been dedicated to using CC as a source of annotated geospatial data. In this paper, we introduce an efficient pipeline to extract annotated user-generated tracks from GPX files found in CC, and the resulting multimodal dataset with 1,416 pairings of human-written descriptions and MultiLineString vector data from the 6 most recent CC releases. The dataset can be used to study people’s outdoor activity patterns, the way people talk about their outdoor experiences, as well as for developing trajectory generation or track annotation models, or for various other problems in place of synthetically generated routes. Our reproducible code is available on GitHub: https://github.com/ilyankou/cc-gpx.
@@ 2.5. Translating descriptions to English
The majority of GPX files we identified are from European websites, and their descriptions are not in English. We use pycld2121212https://github.com/aboSamoor/pycld2 Python library to identify the original language. We remove tracks whose language is identified as Unknown as those often contain majority numbers or incoherent text, as well as all tracks whose descriptions are in languages that have 5 or fewer samples.
@@ 3. Dataset & Potential Applications
The tracks in the resulting dataset represent real-life outdoor activities that were either recorded (i.e., completed) or planned using GIS software. Thus, these routes can be used in place of synthetically generated trajectory datasets, as well as for training trajectory generation ML models that are focused on recreational activities.
The description-MultiLineString pairs can be used to study how people describe their outdoor activities, such as what kinds of landmarks, landscapes, and experiences they choose to mention or ignore, and get an insight into turn-by-turn decisions of users. These pairs can also be used to fine-tune language models to generate human-like descriptions of routes, or to compute location embeddings. Our dataset can become especially powerful when combined with Point of Interest (POI) and/or road segment datasets, such as those from OpenStreetMap [^17] https://welcome.openstreetmap.org/ or Overture [^18] https://overturemaps.org/.
@@ Acknowledgements.
This work was supported by Ordnance Survey & UKRI Engineering and Physical Sciences Research Council [grant no. EP/Y528651/1].